OpenStudy (anonymous):

Someone either jogs the 5 miles to school in 30 min. or rides his bike in 20 min. What is his average speed going to school? Please show me the steps.

OpenStudy (anonymous):

average v=distance/time

OpenStudy (anonymous):

sorry in order to do it with that formula first convert time to seconds and distance to meteres and the velocity is meters/second

OpenStudy (anonymous):

Please can you show me how to do it

OpenStudy (anonymous):

1 meter = 0.000621371192 miles 1 minute=60 seconds so 1800 seconds and 8046.710003073461 meters 8046.720003073461/1800 =4.470400001707478 m/s

OpenStudy (anonymous):

i think it's not velocity. the question is the average speed. .

OpenStudy (anonymous):

velocity is speed but if you need miles per minute you would multiply that velocity by 1609.344/60 because 1 mile is 1609.344 meters and 1 minute is 60 seconds

OpenStudy (anonymous):

just an idea, since the question is the "average". find the velocity/speed given the average of the time, since the distance is constant. jogging and riding a bike ( 30+20)/2, then proceed on solving for the avearge speed. .

OpenStudy (anonymous):

since 20+30=50/2=25, use this for time, since it asked for the average (just an idea)

OpenStudy (anonymous):

answer is 12 mph. but i don't know how to get it

OpenStudy (anonymous):

does m represent for "mile"?

OpenStudy (anonymous):

miles per hr

OpenStudy (anonymous):

okay do what jer said and you will get 5miles/25minutes and then you would convert it to miles/hour by multiplying by 1/60

OpenStudy (anonymous):

actually that wouldnt work you would convert the minutes from 25 minutes to 2.5/6 hours so you would do 5/(2.5/6) which gets you 12

OpenStudy (anonymous):